We aligned FPMI. I also centered ITMX oplev because the light was not hitting on QPD.
Alignment procedure we took was;
1. Align Y arm to the Y end green(Y green trans to PSL is now 195 uW with Y end laser measured temperature 34.14 degC).
2. Aligned IR using PZT2 to Yarm(Now, TRY ~ 0.90).
3. Aligned ITMX monitoring AS spots.
4. Aligned X arm so that TRX maximize.
5. Fine adjusted both BS and X arm(Now, TRX ~ 0.82).
Beam spot position on ETMX looks a little too high & left (from ETMXF camera), but we will leave it until ASS scripts is fixed.